OpenStudy (anonymous):

the integral of (x +2) dx/ x^2 + 4x -9

OpenStudy (anonymous):

Let u=x^2+4x-9 du=2x+4 dx (1/2)du=x+2 dx So your integral becomes: \[\frac{1}{2}\int\limits \frac{du}{u}=\frac{1}{2}\ln|u|+C=\frac{1}{2}\ln|x^2+4x-9|+C\]